# LOGIN: A Large Language Model Consulted Graph Neural Network Training Framework

### v1

@@ 2.3. Homophily Metrics
We utilize the aforementioned homophily metrics to categorize graph datasets and illustrate that our proposed L O G IN framework can help fundamental baselines achieve remarkable performance on graphs exhibiting varying degrees of homophily.
@@ 3.1. Problem Formulation
In this paper, we explore the integration of LLM consultation into GNN training for the node classification task. We follow the transductive setting of node classification: Given some labeled nodes $\mathcal{V_{L}}\subset\mathcal{V}$ in graph $\mathcal{G}=(\mathcal{V},\mathbf{A},\mathbf{X},\left\{s_{n}\right\}_{n\in\mathcal{V}})$, we aim to classify the remaining unlabeled nodes $\mathcal{V_{U}}=\mathcal{V}\backslash\mathcal{V_{L}}$ within the same graph. Formally, the target is to learn a set of GNN parameters $\mathbf{W}$ to predict the unlabeled nodes with the guidance from LLMs:
(3) $f_{\mathbf{W}|\text{LLM}}:(\mathbf{A,X,\left\{s_{n}\right\}_{n\in\mathcal{V}}})\rightarrow\mathbf{Y}.$
Note that the ground truth labels of nodes, the pseudo labels predicted by GNNs, and the pseudo labels from LLMs are denoted as $\mathbf{Y},\mathbf{\hat{Y}},\mathbf{\hat{Y}}^{L}\in\{0,1\}^{N\times C}$ respectively, with $C$ as the number of classes. It is noteworthy that LLMs only involve in training GNNs, and when testing we only use the GNNs trained with LLMs as consultants to predict.
3.2. Overview of LLM Consulted GNN Training
We display the framework of L O G IN on an example graph in Figure 2. In this pipeline, L O G IN seeks guidance from LLMs when GNNs are uncertain about some ambiguous nodes, and feeds responses from LLMs back to retrain GNNs with a complementary coping mechanism. Specifically, there are three key points within this pipeline: how to select uncertain nodes, what the GNN delivers to the LLM, and how the LLM feeds back to the GNN, corresponding to uncertainty measure, LLM consultation, and response feedback.
3.3. Node Selection: Uncertainty Measure
To avoid unnecessary time and resource consumption while aiming to propose a simple yet universal method, we investigate how to identify the crucial timings, i.e., which nodes in this scenario, warrant consultation with LLMs. Regarding this question, we use the variance of the predictions from $\mathbf{T}$ GNNs to represent with different dropouts to represent the predictive uncertainty, with Monte Carlo dropout variational inference (Gal and Ghahramani, 2016) as the mathematical ground, thereby assessing the necessity of consulting LLMs.
Bayesian estimation is used to rate the predictive certainty in common neural networks (Kendall and Gal, 2017). Under the Bayesian framework, the conventionally fixed GNN parameters $\mathbf{W}$ are considered as random variables following specific distributions. The predictive probability of the Bayesian GNN with parameters $\mathbf{W_{b}}$ can be defined as Eq. (4).
@@ 3.3. Node Selection: Uncertainty Measure
Since the true posterior $p(\mathbf{W}\mid\mathbf{A},\mathbf{X})$ in Eq. (4) is hard to calculate in practice, variational inference uses an arbitrary distribution $q_{\theta}(\mathbf{W})$ to approximate the posterior. The key idea of MC dropout variational inference is that dropout could serve to perform variational inference where the variational distribution $q_{\theta}(\mathbf{W})$ is from a Bernoulli variable $\mathbf{M}_{\theta}$, representing whether the neurons are on or off, as Eq. (5) shows.
@@
Note that $\theta$ is the dropout rate, and $\mathbf{M}_{\theta}$ represents a binary mask that controls which neurons in GNNs are off. Then, $\{\hat{\mathbf{W}}_{t}\}_{t=1}^{T}$ is a collection of $\mathbf{T}$ samples drawn from $\mathbf{W}_{\theta}$ in a Monte Carlo way. Through the minimization of the loss function defined in Eq. (6) using these weight samples, $\mathbf{W}_{b}$ is acquired.
@@
3.4. GNN2LLM: LLM Consultation
After we identify the nodes with the most uncertainty $\mathcal{V}_{uc}$ for LLM consultation, the next step in L O G IN is to figure out how to convert the semantic and topological information of these nodes into a format comprehensible by LLMs. In this module, the parameters of LLMs are fixed, and no fine-tuning method is involved. We consult LLMs in a zero-shot node-by-node manner, i.e., constructing a single prompt for each uncertain node, without classification examples provided. Each node prompt is composed of three elements: instruction, input data, and output indicator. The Prompt Construction module in Fig. 2 shows a general prompt example, and we provide the detailed prompts in Appendix A.
@@ 3.4. GNN2LLM: LLM Consultation
The input data includes information related to the target node $n$, namely its original text $s_{n}$, two-hop neighborhood $\mathcal{N}_{2}(n)$ description, and neighbor labels $\mathbf{Y_{\mathcal{N}_{2}(n)}},\mathbf{\hat{Y}_{\mathcal{N}_{2}(n)}}$ from both human annotations and GNN predictions. Instead of listing edges to express connectivity (Chen et al., 2023; Guo et al., 2023) , we opt for summarizing one-hop and two-hop neighbors for the sake of semantic clarity and easier parsing. Note that we select uncertain nodes from the train set, otherwise, it would be a data leakage. The target node’s labels $y_{n},\hat{y}_{n}$ are included in its corresponding prompt to further leverage the sycophancy of LLMs, which is to be addressed in 3.5.2.
3.4.3. Output Parser.
@@
3.5. LLM2GNN: Response Feedback
After consulting LLMs with all selected uncertain nodes $\mathcal{V}_{uc}$, we have a group of LLM predicted pseudo labels and corresponding explanations $\left\{{n\in\mathcal{V}_{uc}}:(\hat{y}_{n}^{L},e_{n})\right\}$. This subsection aims to maximize the utilization of LLM responses and convert the information into signals suitable for GNN processing. We make use of not only correct predictions but also misclassifications made by LLMs. Incorporating such a complementary coping mechanism for response feedback endows GNN training with better interactivity.
Wrong answers from LLMs are confirmed to contain useful knowledge as well (Li et al., 2023). Inspired by this insight, we distinguish between right and wrong LLM responses using ground truth labels and offer distinct approaches to fully leverage the positive and negative data. The right nodes and wrong nodes are represented as $\mathcal{V}_{r}$ and $\mathcal{V}_{w}$ respectively, satisfying $\mathcal{V}_{r}\cup\mathcal{V}_{w}=\mathcal{V}_{uc}$ and $\mathcal{V}_{r}\cap\mathcal{V}_{w}=\emptyset$.
@@ 3.5. LLM2GNN: Response Feedback
For a right node $n\in\mathcal{V}_{r}$, we update its original node embedding $\mathbf{x}_{n}$ with the embedding of the corresponding explanation $e_{n}$ parsed from the LLM response:
(8) $\left\{\boldsymbol{x}_{n}\right\}_{n\in\mathcal{V}_{r}}\leftarrow\text{ENC}\left(\left\{e_{n}\right\}_{n\in\mathcal{V}_{r}})\right.$
@@
Given the emerging reasoning ability that LLMs have shown in natural language tasks (Brown et al., 2020), we opt to attribute the misclassification to the topological information provided in the prompt. We implicitly leverage the sycophancy of LLMs, that the models’ responses are adversely affected by spurious correlations in the input (Weston and Sukhbaatar, 2023), to determine which ego-graph contains misleading links that result in the wrong answer. Around this wrong node $n\in\mathcal{V}_{w}$, we prune edges based on similarity scores to denoise the local structure. As in GNNGuard (Zhang and Zitnik, 2020), we quantify similarity $d_{ni}$ between the wrong node $n$ and its neighbor $i$ using cosine similarity:
@@
(10) $\mathbf{A}\leftarrow\mathbf{1}{\left(\mathbf{D}-d_{th}\cdot\mathbf{I}\right)}\odot\mathbf{M}_{w}\odot\mathbf{A}\$
Note that $\mathbf{D}=\left(d_{ij}\right)_{\mathbf{N}\times\mathbf{N}}$ denotes the similarity matrix, with $\mathbf{M}_{w}$ as a binary node mask representing $\mathcal{V}_{w}$, and $\mathbf{1}$ is a binary indicator:
(11) $\mathbf{1}(d_{ij}-d_{th})=\begin{cases}1&\text{if}~{}{(d_{ij}-d_{th})>0},\\ 0&\text{otherwise}.\end{cases}$
In short, the response feedback stage is divided into two complementary scenarios. When the LLM classifies nodes correctly, we upgrade the original node embeddings with the encoded explanations. Conversely, exploiting the sycophancy of LLMs, we impute the misclassifications to the topological noise, for they might be the spurious correlations provided in prompts, hence pruning edges around the wrong nodes.
@@ 4. Evaluation
In this section, we explore the effectiveness of our L O G IN framework on both homophilic and heterophilic graph datasets, to address the following research questions:
  • RQ1: Does the L O G IN framework achieve performance comparable to state-of-the-art (SOTA) GNNs?
@@ 4.1.1. Datasets
As for heterophilic graphs, we transformed the commonly-used web-page-link graphs: Wisconsin, Texas, and Cornell (Craven et al., 1998) into TAGs by sourcing and incorporating the raw texts [^2] http://www.cs.cmu.edu/~webkb/, which were not available previously in graph libraries. Besides, we fine-tuned DeBERTa-base (He et al., 2021) to encode raw texts into node embeddings. The basic statistics of the datasets are displayed in Table 1. The proportions of heterophilic nodes in graphs are calculated via Definition. 2.2: If the node-level homophily score of node $n$ is below 0.5, then $n$ is defined as a heterophilic node.
@@ 4.1.2. Compared Methods
We compared fundamental GNNs with L O G IN , namely GCN (Kipf and Welling, 2016), GraphSAGE (Hamilton et al., 2017a) and MixHop (Abu-El-Haija et al., 2019), to the advanced state-of-the-art GNNs, including JK-Net (Xu et al., 2018), H2GCN (Zhu et al., 2020), APPNP (Gasteiger et al., 2018), GCNII (Chen et al., 2020), SGC (Wu et al., 2019) ,and SSP (Izadi et al., 2020), in order to demonstrate that the former can achieve performance on par with the latter. MLP also serves as a baseline that predicts without the adjacency information to show the potential out of the simplicity. We select Vicuna-v1.5-7b as our LLM consultant, which is an open-source LLM trained by fine-tuning Llama 2 on user-shared conversations.
For MLP, GCN, GraphSAGE, and MixHop, OR stands for the vanilla backbone model trained with original shallow node features, and FT denotes the vanilla backbone utilizing node embeddings encoded by the fine-tuned LM, while LO represents the model trained within our L O G IN framework. Besides, in the ablation study, we refer to the feature update and structure refinement operations in the response feedback stage as F and S, respectively.
In addition to advanced GNNs, we also compared L O G IN as an implementation of LLMs-as-Consultants paradigm, with the typical methods of LLMs-as-Predictors and LLMs-as-Enhancers, respectively.
@@ 4.1.3. Implementations
We adopt accuracy on the test set of the nodes to evaluate the node prediction performance of all the listed models. And we report the mean accuracy and standard error from five runs with varied random data splits.
For hyper-parameters, in the uncertainty measure module, we implement Monte Carlo dropout variational inference by running models $T$ times with different neurons off, where $T$ represents the number of weight samples in MC dropout. Noted T is always set to 5 in our experiments. For the ratio $\gamma$ of uncertain nodes to consult, we adjust $\gamma$ slightly on each dataset around the proportion of heterophilic nodes shown in Table 1. In the response feedback stage, we tune the similarity threshold $s_{th}$ between $[0.1,0.2]$, according to the off-shell tool GNNGuard (Zhang and Zitnik, 2020).

### v2

@@ 4.3. Homophily Metrics
In this subsection, we introduce homophily metrics defined as follows to distinguish homophilic and heterophilic graphs.
@@
We utilize the aforementioned homophily metrics to categorize graph datasets and illustrate that our proposed L O G IN framework can help classic GNNs achieve remarkable performance on graphs exhibiting varying degrees of homophily.
@@ 5.1. Problem Formulation
In this paper, we explore the integration of LLM consultation into GNN training for the node classification task. We follow the transductive setting of node classification: Given some labeled nodes $\mathcal{V_{L}}\subset\mathcal{V}$ in graph $\mathcal{G}=(\mathcal{V},\mathbf{A},\left\{s_{n}\right\}_{n\in\mathcal{V}})$, we aim to classify the remaining unlabeled nodes $\mathcal{V_{U}}=\mathcal{V}\backslash\mathcal{V_{L}}$ within the same graph. Formally, the target is to learn a set of GNN parameters $\mathbf{W}$ to predict the unlabeled nodes with the guidance from LLMs:
(4) $f_{\mathbf{W}|\text{LLM}}:(\mathbf{A,\left\{s_{n}\right\}_{n\in\mathcal{V}}})\rightarrow\mathbf{Y}.$
Note that the ground truth labels of nodes, the pseudo labels predicted by GNNs, and the pseudo labels from LLMs are denoted as $\mathbf{Y},\mathbf{\hat{Y}},\mathbf{\hat{Y}}^{L}\in\{0,1\}^{N\times C}$ respectively, with $C$ as the number of classes.
5.2. Overview of L O G IN
Figure 2 demonstrates the framework of L O G IN . In this pipeline, we first identify target nodes that need consultation with LLM based on the GNN’s predictive uncertainty (c.f. Section 5.3). Then prompts consisting of the associated texts and local structures of these nodes are provided to LLMs for consultation (c.f. Section 5.4). After LLM consultation, a complementary mechanism is applied to fully utilize the LLMs’ responses, regardless of their classification correctness (c.f. Section 5.5).
5.3. Node Selection Based on Uncertainty
An intuitive node selection approach is to consult the LLM with all the nodes on the graph. However, such a manner fails to utilize the capacity of GNNs, which are proficient at handling simple nodes, and it could be inefficient due to the limited interaction speed with LLMs. Hence, we decide to consult LLMs only with partial difficult nodes in the graph, and adopt the variance of the predictions from several GNNs as a surrogate indicator for identifying the hard ones in our L O G IN .
Recall that Bayesian estimation is typically used to rate the predictive certainty in common neural networks (Kendall and Gal, 2017). Under the Bayesian framework, the conventionally fixed GNN parameters $\mathbf{W}$ are considered as random variables following specific distributions. The predictive probability of the Bayesian GNN with parameters $\mathbf{W_{b}}$ can be defined as Eq. (5).
@@ 5.3. Node Selection Based on Uncertainty
Since the true posterior $p(\mathbf{W}\mid\mathbf{A},\mathbf{X})$ in Eq. (5) is hard to calculate in practice, variational inference uses an arbitrary distribution $q_{\theta}(\mathbf{W})$ to approximate the posterior. Specifically, we use MC dropout variational inference, in which dropout could serve to perform variational inference. In this method, the variational distribution $q_{\theta}(\mathbf{W})$ is from a Bernoulli variable $\mathbf{M}_{\theta}$, representing whether the neurons are on or off, as Eq. (6) shows.
@@
where $\theta$ is the dropout rate, and $\mathbf{M}_{\theta}$ represents a binary mask that controls which neurons in GNNs are off. Then, $\{\hat{\mathbf{W}}_{t}\}_{t=1}^{T}$ is a collection of $T$ samples drawn from $\mathbf{W}_{\theta}$ in a Monte Carlo way. Through the minimization of the loss function defined in Eq. (7) using these weight samples, $\mathbf{W}_{b}$ can be acquired.
@@
5.4. LLM Consultation
After we identify the nodes with the most uncertainty $\mathcal{V}_{uc}$ for LLM consultation, the next step in L O G IN is to figure out how to convert the semantic and topological information of these nodes into an understandable format by LLMs. We consult LLMs in a zero-shot node-by-node manner, i.e., constructing a single prompt for each uncertain node, without classification examples provided. Through this way, no fine-tuning method is involved, and the parameters of LLMs are fixed. Each node prompt is composed of three elements: instruction, input data, and output indicator. The Prompt Construction module in Fig. 2 shows a general prompt example, and we provide the detailed prompts in Appendix A.
@@ 5.4. LLM Consultation
The input data includes information related to the target node $n$, namely its original text $s_{n}$, the two-hop neighborhood $\mathcal{N}_{2}(n)$ description, and the neighbor labels $\mathbf{Y_{\mathcal{N}_{2}(n)}},\mathbf{\hat{Y}_{\mathcal{N}_{2}(n)}}$ from both human annotations and GNN predictions. Instead of listing edges to express connectivity (Chen et al., 2023; Guo et al., 2023) , we opt to summarize one-hop and two-hop neighbors for the sake of semantic clarity and easier parsing. Note that we only select uncertain nodes from the train set, otherwise, it might bring a data leakage. The target node’s labels $y_{n},\hat{y}_{n}$ are also included in its corresponding prompt to provide more information for LLM predictions.
5.4.3. Output Indicator.
@@
5.5. LLM Response Feedback
After consulting LLMs with all selected uncertain nodes $\mathcal{V}_{uc}$, we have a group of LLM predicted pseudo labels and corresponding explanations $\left\{{n\in\mathcal{V}_{uc}}:(\hat{y}_{n}^{L},e_{n})\right\}$. Next, we aim to maximize the utilization of LLM responses and convert the information into signals suitable for GNN processing. The LLM responses can be divided into two types compared to ground truth labels, i.e., the correct predictions when $\hat{y}_{n}^{L}=y_{n}$ and the wrong predictions when $\hat{y}_{n}^{L}\neq y_{n}$. We not only utilize the correct predictions but also the misclassifications, and offer distinct approaches respectively. Incorporating such a complementary coping mechanism endows GNN training with full exploitation of LLM responses.
For clarity in expression, we denote the correct nodes and incorrect nodes as $\mathcal{V}_{r}$ and $\mathcal{V}_{w}$ respectively, satisfying $\mathcal{V}_{r}\cup\mathcal{V}_{w}=\mathcal{V}_{uc}$ and $\mathcal{V}_{r}\cap\mathcal{V}_{w}=\emptyset$.
@@ 5.5. LLM Response Feedback
For a right node $n\in\mathcal{V}_{r}$, we append the explanation $e_{n}$ parsed from the LLM response to its original text $s_{n}$. And we attain its new node embedding $\mathbf{{x}_{n}}^{\prime}$ by encoding the new attached text from $e_{n}$:
(9) $\left\{\mathbf{{x}_{n}}^{\prime}\right\}_{n\in\mathcal{V}_{r}}\leftarrow\text{ENC}\left(\left\{s_{n}+e_{n}\right\}_{n\in\mathcal{V}_{r}})\right.$
@@
For those wrong nodes, the LLM responses are left out since the explanations for incorrect classifications could be unhelpful. Instead, we opt to attribute the misclassification to the topological information, e.g., complex structure patterns and potential adversarial noises. Therefore, we heuristically decide to simplify the local structure for the wrong nodes to make it easier for GNNs to learn. Specifically, around a wrong node $n\in\mathcal{V}_{w}$, we prune edges based on node similarity scores to denoise the local structure. As in representative graph structure learning methods, e.g. GNNGuard (Zhang and Zitnik, 2020), we measure similarity $d_{ni}$ between the features of the wrong node $n$ and its neighbor $i$ using cosine similarity:
@@
(11) $\mathbf{A^{\prime}}\leftarrow\mathbf{1}{\left(\mathbf{D}-d_{th}\cdot\mathbf{I}\right)}\odot\mathbf{M}_{w}\odot\mathbf{A}\$
Here $\mathbf{D}=\left(d_{ij}\right)_{\mathbf{N}\times\mathbf{N}}$ denotes the similarity matrix, with $\mathbf{M}_{w}$ as a binary node mask representing $\mathcal{V}_{w}$, and $\mathbf{1}$ is a binary indicator.
In a nutshell, the response feedback stage is divided into two complementary scenarios. When the LLM classifies nodes correctly, we upgrade the original node embeddings with the encoded explanations. Conversely, we impute the misclassifications to the topological noise, hence pruning edges around the wrong nodes.
5.6. Train and Test
After the LLM response feedback stage, L O G IN is set to retrain GNN. When retraining, the GNN aggregates the original node features and the ones updated with the correct classification grounds. And for the nodes that LLMs misclassify, aggregation only happens along the retained edges after structure refinement. In this way, the message passing of GNNs is integrated with the LLM intelligence both semantically and topologically.
The GNN training process with one-time consultation within the L O G IN framework is illustrated in Algorithm 1. This process can be iteratively repeated until the maximum number of iterations is reached or the performance of the GNN has achieved an acceptable level.
During the test, the ground truth labels are inaccessible, so we only use the trained GNN without additional LLM consultation to make final predictions. Since we retrain the GNN over the graphs augmented by LLM consultation, i.e., with the node representations enhanced and the noisy links removed, the attained GNNs can perform aggregation in a more reliable neighborhood with more accurate semantic information and thus produce improvements.
Input: $\mathcal{G}=(\mathcal{V},\mathbf{A},,\left\{s_{n}\right\}_{n\in\mathcal{V}})$: A TAG, $\mathcal{V}_{\mathrm{train}}$: Set of training nodes, $\mathbf{X}$: Node representations, $\mathbf{Y}$: Labels of training nodes, $T$: Number of Monte Carlo samples, $\theta$: Dropout rate for uncertainty estimation, $\gamma$: Ratio of uncertain nodes, $d_{th}$: Similarity threshold
Output: The retrained GNN parameters after LLM consultation $\mathbf{W}_{r}$.
1
2Choose basic models for pretrained GNN $f_{\mathbf{W}_{p}}$ and retrained GNN $f_{\mathbf{W}_{r}}$;
3
4Initialize parameters $\mathbf{W}_{p}$ and $\mathbf{W}_{r}$;
5
6for $t=1,\ldots,T$ do
7       Sample $\hat{\mathbf{W}}_{t}$ from $\mathbf{W_{\theta}}$
8Train $\mathbf{W}_{p}$ and $\{\hat{\mathbf{W}}_{t}\}_{t=1}^{T}$ by minimizing Eq. (7);
9 Estimate the model uncertainty $U$ w.r.t. Eq. (8);
10 Divide $\mathcal{V}_{\mathrm{train}}$ into $\mathcal{V}_{\mathrm{c}}$ and $\mathcal{V}_{\mathrm{uc}}$ according to $U$ such that $\gamma=\frac{|\mathcal{V}_{\mathrm{uc}}|}{|\mathcal{V}_{\mathrm{train}}|}$;
11 for $n\in\mathcal{V}_{\mathrm{uc}}$ do
12       Consult the LLM with prompt $n$ and gain LLM-predicted label and its explanation $\hat{y_{n}}^{L},e_{n}$ ;
13
14      if $\hat{y_{n}}^{L}=y_{n}$ then
15             Update node features $\boldsymbol{x}_{n}$ w.r.t. Eq. (9);
16      else
17            Prune edges of $n$ w.r.t. Eq. (11);
18
19Retrain GNN $f_{\mathbf{W}_{r}}$ with new $\mathbf{A^{\prime}},\mathbf{X^{\prime}}$;
Return $\mathbf{W}_{r}$.
Algorithm 1 L
O
G
IN
@@ 6. Evaluation
In this section, we investigate the effectiveness of our L O G IN framework on both homophilic and heterophilic graph datasets, to address the following research questions:
  • RQ1: Does the L O G IN framework achieve performance comparable to state-of-the-art GNNs?
@@
  • RQ6: Can consulting LLMs with more training nodes achieve higher performance increases with classic GNNs?
  • RQ7: Can L O G IN also help advanced GNNs enhance performance?
@@ 6.1.1. Datasets
For heterophilic graphs, we transformed the commonly-used web-page-link graphs: Wisconsin, Texas, and Cornell (Craven et al., 1998) into TAGs by sourcing and incorporating the raw texts [^1] http://www.cs.cmu.edu/~webkb/, which were not available previously in graph libraries. Besides, we fine-tuned DeBERTa-base (He et al., 2021) to encode raw texts into node embeddings. The basic statistics of the datasets are displayed in Table 1. The proportions of heterophilic nodes in graphs are calculated via Definition 4.2: If the node-level homophily score of node $n$ is below 0.5, then $n$ is defined as a heterophilic node.
@@ 6.1.2. Compared Methods
We integrate fundamental GNNs with L O G IN , namely GCN (Kipf and Welling, 2016), GraphSAGE (Hamilton et al., 2017a) and MixHop (Abu-El-Haija et al., 2019), to compare with the advanced state-of-the-art GNNs, including JK-Net (Xu et al., 2018), H2GCN (Zhu et al., 2020), APPNP (Gasteiger et al., 2018), GCNII (Chen et al., 2020), SGC (Wu et al., 2019) ,and SSP (Izadi et al., 2020), to demonstrate that the former can achieve performance on par with the latter. MLP, which predicts without adjacency information, also serves as a base learner combined with L O G IN to show the potential of our framework. We select Vicuna-v1.5-7b as our LLM consultant, an open-source LLM trained by fine-tuning Llama 2 on user-shared conversations.
For MLP, GCN, GraphSAGE, and MixHop, “OR” stands for the vanilla backbone model trained with original shallow node features, and “FT” denotes the vanilla backbone utilizing node embeddings encoded by the fine-tuned LM. “LO” represents the model trained within our L O G IN framework. Besides, in the ablation study, we refer to the feature update and structure refinement operations in the LLM response feedback stage as “F” and “S”, respectively.
In addition to advanced GNNs, we also compared L O G IN as an implementation of LLMs-as-Consultants paradigm, with the other existing paradigms, i.e. LLMs-as-Predictors and LLMs-as-Enhancers, respectively.
@@ 6.1.3. Implementations
We adopt accuracy on the test set of the nodes to evaluate the node prediction performance of all the listed models. We report the mean accuracy and standard error from five runs with varied random data splits, and all the reported results are statistically significant.
For hyper-parameters, in the uncertainty measure module, we implement Monte Carlo dropout variational inference by running models $T$ times with different neurons off, where $T$ represents the number of weight samples in MC dropout. $T$ is always set to 5 in our experiments. For the ratio $\gamma$ of uncertain nodes to consult, we adjust $\gamma$ slightly on each dataset around the proportion of heterophilic nodes shown in Table 1. In the response feedback stage, we tune the similarity threshold $s_{th}$ between $[0.1,0.2]$, according to the off-shell tool GNNGuard (Zhang and Zitnik, 2020).